jarod in having a problem with rabbits in his veggie garden so he decides to fence it in. the length of the garden is 7 ft more than 3 times the width he needs 78 ft of fencing to do the job.
Answer by rapaljer(4671)

Let x = width
3x+7= length

2( ) + 2( )= Perimeter
2(x) +2(3x+7)=78
2x+6x+14=78
8x=78-14
8x=64
x=64/8=8 feet = width

3x+7= length
3(8)+7 = length
24+7=31 feet

Check: 2*8+2*31
16+ 62 =78 so it checks!!

For additional help with square roots, please see my own website. Click on my tutor name "rapaljer" anywhere in algebra.com. There, click on my website. From my Homepage, look for the link near the top of the page "Basic, Intermediate, and College Algebra: One Step at a Time." Then click on "Basic Algebra", then "Chapter 1", "Sections 1.09 and 1.10." There you will find my own non-traditional, step by step explanations, examples, exercises, and answers to ALL the exercises, and solutions to many of the problems on the "MATH IN LIVING COLOR" pages that go with these explanations. I have LOTS of exercises that are solved for you in a way that will be a lot easier to understand than most of the expensive textbooks of today. And, like algebra.com, it's all FREE!!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus